It has long been stated that the mean temperature of humans is 98.6 degrees°F.​However, two researchers currently involved in the subject thought that the mean temperature of humans is less than 98.6 degrees°F.They measured the temperatures of 50 healthy adults 1 to 4 times daily for 3​ days, obtaining 225 measurements. The sample data resulted in a sample mean of

98.3 degrees°F and a sample standard deviation of 1.1 degrees°F. Use the​ P-value approach to conduct a hypothesis test to judge whether the mean temperature of humans is less than

98.6 degrees°F at the alpha α equals=0.01 level of significance.

I need to find the P-value

I need to know if it will be rejected or not.

I am having trouble using the T-Table..
Answer #1

Given that, sample size (n) = 225

sample mean (\bar x) = 98.3 °F and

sample standard deviation (s) = 1.1 °F

The null and alternative hypotheses are,

H0 : μ = 98.6

Ha : μ < 98.6

Since, sample size is very large we used standard normal distribution (z-test).

Test statistic is,

- 98.3 - 98.6 -= 225 -4.09

=> Test statistic is, Z = -4.09

p-value = P(Z < -4.09) = 0.0000

=> p-value = 0.0000

Since, p-value <  α = 0.01, we reject the null hypothesis (H0).

Note : In t-table df = 225 - 1 = 224 is not given so, we need to used Excel or TI-83 or TI-84 calculator to find p-value.

Excel Command : =TDIST(4.09, 224, 1) = 0.0000

​​​

​​​​​